OpenStudy (anonymous):

Find the distance between the skew lines L1: x=y=z and L2: x+1=1/2y=1/3z

OpenStudy (amistre64):

we need a line that is perp to both of them, so cross their direction vectors is my first idea

OpenStudy (anonymous):

what would be the parametric equation for line2?

OpenStudy (amistre64):

these are formated such that t = .... line 1 x = y= z = t x = 0+t y = 0+t z = 0+t line2 x+1=1/2y=1/3z = t x+1 = t ; x = -1+t 1/2 y = t ; y = 0+2t 1/3 z = t ; z = 0+3t

OpenStudy (anonymous):

thank you . i got it from here

OpenStudy (amistre64):

youre welcome